OpenStudy (geekfromthefutur):

I give medal and fan also testimonial A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (jadeishere):

minimum is 0.425. And the maximum is 0.625

OpenStudy (geekfromthefutur):

Well how did u get that lol

OpenStudy (jadeishere):

lol hold on i need to email my teacher an assignment real quick

OpenStudy (jadeishere):

wait i did it wrong >.<

OpenStudy (jadeishere):

Sowwy it's basic addition and subtraction. That's where i went wrong the first time >.< 0.525 - 10 = 0.515 and 0.525 + 10 = 0.535

OpenStudy (anonymous):

if there is a tolerance that means the drill bit can be -slightly above or -slightly below soo The first equation is x - 0.0525 = 0.01. The 2nd is: x - 0.0525 = -0.01 so the answer to the first one is 0.0625 so the second one is x=0.0425 0.0425 - 0.0525 = -0.01, |-0.01|=0.01 So the answer to the second one is 0.0425 So the minimum diameter of the drill bits is 0.425. And the maximum is 0.625. both have a deviation of no more than 0.01

OpenStudy (jadeishere):

^you catch my mistake?

OpenStudy (anonymous):

oh nvm >.< lol...>.>

OpenStudy (jadeishere):

Sorry i felt like trying to see if you'd catch that 0.01 is 100 not 10 xD